| <p>On 25/11/2018:</p>
<p>The beam was tuned for 6.0MeV 7Be 2+</p>
<p>Using 10mm aperture FCC = 12 pA</p>
<p>Run # 181125121226</p>
<ul>
<li>Run time: 7264s</li>
<li>Pulser ON, as before</li>
<li>Target: CH2 XVIII</li>
<li>Beam: 6.0MeV 7Be 2+</li>
<li>See attached picture 1 for sample spectrum of Far S2 front strip. See attached picture 2 for diode spectrum.</li>
</ul>
<p>Run # 181125141350</p>
<ul>
<li>Run time: 1887s</li>
<li>Pulser ON, as before</li>
<li>Target: CH2 XVIII</li>
<li>Beam: 6.0MeV 7Be 2+</li>
</ul>
<p>During this run at 14:40 there was a very brief power outage. The run was stopped afterwards.</p>
<p>Check chamber setup:</p>
<p>Far S2 bias = -130.0V, leak I = 1.89uA</p>
<p>Diode bias = -70.1V, leak I = 0.00uA</p>
<p>Pressures (mBar):</p>
<ul>
<li>TPU1 = 1.0e-6</li>
<li>TPU2 = 3.2e-7</li>
<li>TPD1 = 6.5e-7</li>
<li>TPD2 = 6.1e-7</li>
<li>CHUP = Sensor Off</li>
<li>CHDW = No Sensor</li>
</ul>
<p>Preamps for Far S2 and diode are still drawing correct currents. The chamber setup appears unaffected by the brief outage.</p>
<p>At this point the pulser was disconnected from the oscilloscope - removing the load on the pulser. We expect the pulser peak position to change as a result.</p>
<p>After the outage we needed to retune beam for 6.0MeV 7Be 2+</p>
<p>Moved target ladder to position #2 - blank 10mm aperture. Had troubles finding beam, moved target to natC III. In diode spectrum we saw beam energies ~ 7-8MeV.</p>
<p>It turned out the bending magnet located right after the ion splutter source needed recalibrating. It was recalibrated using Hydrogen (A=1), Carbon (A=12), and Oxygen (A=16).</p>
<p>During the retuning, the pulser frequency was checked for both the 2.5 and 25Hz settings:</p>
<p>Pulser setting set to 2.5Hz</p>
<ul>
<li>Pulser frequency measured internally by scope: 2.095 -> 2.096Hz</li>
<li>Pulser frequency measured using scopes cursors: 2.092Hz</li>
</ul>
<p>Pulser setting set back to 25Hz</p>
<ul>
<li>Pulser frequency measured internally by scope: 19.29 -> 19.30Hz</li>
<li>Pulser frequency measured using scopes cursors: 19.38Hz</li>
</ul>
<p>After recalibrating the bending magnet, we switched beam to 6.0MeV 7Li 2+ to benefit from the higher beam current.</p>
<p>After tuning: FC4 = 1.25nA. Using blank 10mm aperture FCC = 96pA, using blank 3mm aperture FCC = 66pA -> ~70% transmission attained.</p>
<p>Switched beam back to 6.0MeV 7Be 2+</p>
<p>Using 3mm aperture FCC = 7pA.</p>
<p>Opened Far S2 iris (was closed during retuning after power outage), set target ladder to CH2 XVIII</p>
<p>Run # 181125223800</p>
<ul>
<li>Run time: 35914s</li>
<li>Pulser ON, 25Hz</li>
<li>Target: CH2 XVIII</li>
<li>Beam: 6.0MeV 7Be 2+</li>
<li>Left running overnight</li>
<li>See attached picture 3 for sample spectrum of Far S2 front strip. See attached picture 4 for diode spectrum.</li>
</ul> |